FOURIER-TRANSFORM SPECTROSCOPY OF I_{2} (D-X) EMISSION FOLLOWING OPTICAL-OPTICAL DOUBLE RESONANCE EXCITATION OF THE I_{2} (E) STATE

\underline{VADIM A. ALEKSEEV}, Institute of Physics, St.Petersburg State University, Peterhof, 198504 Russia; AMANDA J. ROSS, Laboratoire de Spectrométrie Ionique et Moléculaire, Université’ de Lyon (Lyon 1), 69622 Villeurbanne, France.

The E 0_{g}^{+} ion-pair state of I_{2} molecule may be excited by the optical-optical double resonance (OODR) via the valence state B 0_{u}^{+}, I_{2}(X 0_{g}^{+}) + hv_{1} → I_{2} (B 0_{u}^{+} ) + hv_{2} → I_{2} ( E 0_{g}^{+} ). Emission spectra of pure I_{2} vapor selectively excited in this way show not only ransitions from the E 0_{g}^{+} state, but also D 0_{u}^{+} → X 0_{g}^{+} transition in the near ultraviolet. Typically the populated levels in the D 0_{u}^{+} state are located less than 300 cm^{−1} below the laser-excited level. In this contribution we examine rotationally resolved Fourier-transform spectra of the D 0_{u}^{+} → X 0_{g}^{+} emission following OODR excitation of various E 0_{g}^{+} v, J levels with single mode cw dye lasers. We show that the E 0_{g}^{+} → D 0_{u}^{+} population transfer obeys the J_{D}=J_{E} _{−}^{+} 1, 3, 5 ... propensity rule. Transfer to the J_{D}=J_{E} _{−}^{+}1 levels is the most probable although in some cases, J_{D}=J_{E}-1 and J_{D}=J_{E}+1 have significantly different populations.